Find the coordinates of the points of intersection of the line with equation y=x+k and the parabola with equation y=x^2-4x, where k>0
Answer by Natolino1983(23)

Y=(x+k)=x^2-4x then, x^2-5x-k=0 must be solved.
Formula says x=(5+sqrt(25 +4k))/2 or (5-sqrt(25+4k))/2.
The points are P=((5+sqrt(25+4k))/2,(5+sqrt(25+4k)+2k)/2) and Q=((5-sqrt(25-4k))/2,5-sqrt(25+4k)+2k)/2)